### v1

@@ 3.2 Black Hole Spins
While there are still large uncertainties in the angular momentum transport mechanisms during the gravitational collapse of heavy stars and the formation of black holes, most models suggest that the formation of fast-rotating black holes from stellar collapses are unlikely Woosley:2006fn ; Yoon:2006fr . Moreover, it is often thought that the value of the spin parameter of stellar-mass black holes is natal King:1999aq ; Valsecchi:2010cw ; Wong:2011eg . If the black hole is in a low-mass X-ray binary, the black hole cannot significantly change its spin parameter $a_{*}$ even swallowing the whole companion star, because $M_{\rm BH}\sim 10$ $M_{\odot}$ and $M_{\rm star}\lesssim M_{\odot}$. If the black hole is in a high-mass X-ray binary, the lifetime of the companion star is too short and it is impossible to transfer enough material from the companion star to the black hole even if the latter accretes at its Eddington limit. However, alternative models also exist. Fragos & McClintock (2015) Fragos:2014cva proposed that a black hole in a low-mass X-ray binary can be spun up immediately after its formation. Qin et al. (2019) Qin:2018sxk proposed that the black hole spin in a high-mass X-ray binary may be directly related to the angular momentum of the progenitor star, which can transfer part of its envelope to the companion star and get a core with high angular momentum.
@@
Low-mass X-ray binaries are not progenitors of binary black holes because the companion stars have masses $M_{\rm star}\lesssim M_{\odot}$ and therefore they evolve into white dwarfs, not black holes. High-mass X-ray binaries can be progenitors of binary black holes, but the observed high-mass black hole X-ray binaries formed at high metallicity: the companion stars are massive, so their lifetime is short and they have to be formed recently, which is also confirmed by the fact the masses of the black holes are $M_{\rm BH}\lesssim 20$ $M_{\odot}$. On the other hand, it is also true that spin measurements of black holes in X-ray binaries are affected by selection effects and systematic uncertainties. Robust spin measurements are possible for fast-rotating black holes with compact coronae, while it is difficult to distinguish fast-rotating black holes with extended coronae and slow-rotating black holes with compact coronae Dauser:2013xv ; Shashank:2025hka .
@@ 3.3 Is General Relativity Correct?
With gravitational wave data, we can test specific gravity models Yunes:2016jcc . The gravitational wave signal does not depend on the spacetime geometry only, but even on how matter/gravity curves the spacetime, how gravitational waves are produced, and on the properties of the gravitational waves. It is not possible to test the Kerr metric in a model independent way as in the case of electromagnetic tests. Alternatively, we can look for deviations from the gravitational wave signal expected in General Relativity or we can perform a number of consistency tests LIGOScientific:2016lio ; LIGOScientific:2019fpa ; LIGOScientific:2020tif ; LIGOScientific:2021sio . See Ref. Colleoni:2024lpj for a recent review on tests of General Relativity with gravitational wave data of the coalescence of stellar-mass black holes.
@@ 5 Concluding Remarks
This work was supported by the National Natural Science Foundation of China (NSFC), Grant No. 12261131497 and No. 12250610185.
[…]
No new data were created in this work.
[…]
The author declares no conflicts of interest.
{adjustwidth}
[…]

### v2

@@ 3.2 Black Hole Spins
While there are still large uncertainties in the angular momentum transport mechanisms during the gravitational collapse of heavy stars and the formation of black holes, most models suggest that the formation of fast-rotating black holes from stellar collapses are unlikely Woosley:2006fn ; Yoon:2006fr . Moreover, it is often thought that the value of the spin parameter of stellar-mass black holes is natal King:1999aq ; Valsecchi:2010cw ; Wong:2011eg . If the black hole is in a low-mass X-ray binary, the black hole cannot significantly change its spin parameter $a_{*}$ even swallowing the whole companion star, because $_{\rm BH}\sim 10 MM_{\odot}$ and $M_{\rm star}\lesssim M_{\odot}$. If the black hole is in a high-mass X-ray binary, the lifetime of the companion star is too short and it is impossible to transfer enough material from the companion star to the black hole even if the latter accretes at its Eddington limit. However, alternative models also exist. Fragos & McClintock (2015) Fragos:2014cva proposed that a black hole in a low-mass X-ray binary can be spun up immediately after its formation. Qin et al. (2019) Qin:2018sxk proposed that the black hole spin in a high-mass X-ray binary may be directly related to the angular momentum of the progenitor star, which can transfer part of its envelope to the companion star and get a core with high angular momentum.
@@
Low-mass X-ray binaries are not progenitors of binary black holes because the companion stars have masses $M_{\rm star}\lesssim M_{\odot}$ and therefore they evolve into white dwarfs, not black holes. High-mass X-ray binaries can be progenitors of binary black holes, but the observed high-mass black hole X-ray binaries formed at high metallicity: the companion stars are massive, so their lifetime is short and they have to be formed recently, which is also confirmed by the fact the masses of the black holes are $_{\rm BH}\lesssim 20 MM_{\odot}$. On the other hand, it is also true that spin measurements of black holes in X-ray binaries are affected by selection effects and systematic uncertainties. Robust spin measurements are possible for fast-rotating black holes with compact coronae, while it is difficult to distinguish fast-rotating black holes with extended coronae and slow-rotating black holes with compact coronae Dauser:2013xv ; Shashank:2025hka .
@@ 3.3 Is General Relativity Correct?
With gravitational wave data, we can test specific gravity models Yunes:2016jcc . The gravitational wave signal does not depend on the spacetime geometry only, but even on how matter/gravity curves the spacetime, how gravitational waves are produced, and on the properties of the gravitational waves. It is not possible to test the Kerr metric in a model independent way as in the case of electromagnetic tests. Alternatively, we can look for deviations from the gravitational wave signal expected in General Relativity or we can perform a number of consistency tests LIGOScientific:2016lio ; LIGOScientific:2019fpa ; LIGOScientific:2020tif ; LIGOScientific:2021sio ; Das:2025riq . See Ref. Colleoni:2024lpj for a recent review on tests of General Relativity with gravitational wave data of the coalescence of stellar-mass black holes.